OpenStudy (anonymous):

prove that the line y=3x-10 does not intersect the curve y=x^2-5x+7

OpenStudy (anonymous):

putting y=x^2-5x+7 in the first equation we get x^2-5x+7=3x-10 x^2-8x+17=0 in the quadratic equation discriminator D=8^2-4*17=-10 so roots are imaginary hence the curve does not intersect